OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = sin 5x + sin^2 5x, (0, 0) I got 5x but it was wrong please help me

OpenStudy (loser66):

y'= ??

OpenStudy (anonymous):

cos5x+2sin5x?

OpenStudy (anonymous):

Forgot chain rule.

OpenStudy (loser66):

it's not that but I got what you have. hahahaha....

OpenStudy (anonymous):

\(y' = cos(5x)\cdot5 +2sin(5x)\cdot cos(5x)\cdot5\) \(y' = 5cos(5x) + 10sin(5x)cos(5x)\)

OpenStudy (loser66):

@Concentrationalizing some more steps, you will get what he got and it is wrong (according to him) :)

OpenStudy (loser66):

but I don't know how it is wrong!! or you should write y = 5x @gaba

OpenStudy (anonymous):

Plug in 0 into the derivative: \(y' = 5cos(0) + 10sin(0)cos(0)\) \(y' = 5\) Hmm....question is typed incorrectly?

OpenStudy (anonymous):

i needed to type it as y=5x! thank you!!

OpenStudy (anonymous):

oh, lol. Okay, cool.

OpenStudy (loser66):

hahahaha...................... it's fun.